The electric motor of a model train accelerates the train from rest to 0.560 m/s in 31.0 ms. The total mass of the train is 505 g. Find the average power delivered to the train during its acceleration.
Answer #1

The acceleration a of the train is ∆v/∆t = (0.560 m/s)/(0.0310 s) = 18.06451612903226 m/s2
.The force on the train is therefore F = ma = (0.505 kg)*18.06451612903226 m/s2
= 9.12258064516129 N.

During this time is travels a distance given by

x = Vavg*t

= 0.5 × 0.560m/s × 0.031 s

= 0.00868 m.

The average power is  P = F x/t = 9.12258064516129 N × 0.00868 m/0.031 s

= 2.554322580645161 W